if a= 2+3 1/2 find a 1/2 +1/a 1/2

 
if a= 2+31/2
 
find a1/2 +1/a1/2

Grade:10

3 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
Given that a=2+\sqrt{3}
So\sqrt{a}+1/\sqrt{a}=(a+1)/\sqrt{a}=(3+\sqrt{3})/\sqrt{2+\sqrt{3}}
This value should be calculated directly and it cannot be reduced further
after calculating we get its value as 2.45
Yugabrat Gogoi
9 years ago
Please find your answer below.
 
According to me the answer should be as follows.
 
a1/2+ 1/a1/2
=(2+31/2)+1/(2+31/2)1/2
={ [underroot 2+(underroot3)]2+1}/underroot2+(underroot3)
=3+root3/underroot 2+(underroot3)
 
 
if there’s a mistake please tell me. regards from yugabrat gogoi
Yugabrat Gogoi
9 years ago
the approx value is2.449489743.